OpenStudy (anonymous):

[(x-3)+y] [(x-3)+y] I got two answers for these problems which one is correct? x^2 - 6x +y^2 and x^2 -6x +9 + 2xy + y^2

OpenStudy (shayaan_mustafa):

both will be correct. coz it seems like quadratic equation dear.

OpenStudy (anonymous):

\[\Large [(x-3)+y]\cdot [(x-3)+y]=(x-3+y)\cdot (x-3+y)\] \[\Large x^2-3x+xy-3x+9-3y+xy-3y+y^2\] \[\Large x^2-6x-6y+2xy+9+y^2\]

OpenStudy (shayaan_mustafa):

@Kreshnik how you magnify your equations? when you write equation then it becomes large font. how?

OpenStudy (anonymous):

\[\huge x^2-4+23...\] --- this one was \huge then write equation... I write mostly with .. \Large ... BUT sometimes \LARGE THE ONE YOU ASKED WAS TYPED WITH \Large :)

OpenStudy (anonymous):

neither of your answer is correct. @Kreshnik's answer is the right answer.

OpenStudy (shayaan_mustafa):

i don't understand. can you kindly make a syntax for this?

OpenStudy (anonymous):

syntax means ?? :( @Shayaan_Mustafa ... (Do you know to use Latex , down you have Equation button see??)

OpenStudy (shayaan_mustafa):

yes. i can see. i know latex.

OpenStudy (anonymous):

.... Here's an example ... \Large x^2-2+4 \cdot 0 Copy this and click Equation button... then PASTE , Enjoy. ;)

OpenStudy (shayaan_mustafa):

\[\huge x ^{2}\]

OpenStudy (anonymous):

\[\Large x^2-2+4 \cdot 0\]

OpenStudy (shayaan_mustafa):

aha. thnx. thnx a lot. i got it.

OpenStudy (anonymous):

@Shayaan_Mustafa Well done !!